How many yards of 50" fabric is needed to cover a round table with a 32" surfice with a 5" over hang.
Answer by stanbon(75887)

How many yards of 50" fabric is needed to cover a round table with a 32" surfice with a 5" over hang.
----
diameter = 50" + 2*5" = 60"
-----
Area of fabric = pi*radius^2 = pi*30^2 = 900pi sq inches
Each sq yrd is 36*36 = 1296 sq in
------
Ans: (900pi/1296) = 2.18 sq yd
----------------
Cheers,
Stan H.
---------------